The frequency of yellow light is
4.76 × 1014 Hz. Find its wavelength. The speed of light is
2.99792 × 108 m/s . Answer in units of m
1
Expert's answer
2013-03-12T13:16:39-0400
λ=с/f
where λ -wavelength, c - speed of light, f - frequency so, λ=2.99792*10^8/(4.76*10^14)=630*10^-9 (m) =
6.3*10^-7 (m) or 630 (nm)But it is not a wavelegth of yellow light, it is red.